Fine 2 Part A: Standardization of EDTA Concentration of CaCl, solution ().01 m Rough Fine 1 Volume of CaCl, Initial Volume ED EDTA concentrations for each titration
Answer #1

Calcium form complex with EDTA in 1:1 stoichiometric ratio.

So, using M1V1 (CaCl2) = M2V2 (EDTA)

For fine 1 : 0.01×10 = M2×9.44

EDTA concentration, M2 = 0.1/9.44 = 0.01 M

For fine 2 : 0.01×10 = M2 ×9.38

EDTA concentration = 0.1/9.38 = 0.01 M

Average concentration of EDTA = (0.01+0.01)/2 = 0.01 M

Please give feedback!